OpenStudy (anonymous):

7y-(9y-15)=14-(-y+19)

jimthompson5910 (jim_thompson5910):

7y-(9y-15)=14-(-y+19) 7y-9y+15=14+y-19 -2y+15=14+y-19 -2y+15=y-5 -2y=y-5-15 -2y-y=-5-15 -3y=-5-15 -3y=-20 y=-20/(-3) y=20/3

OpenStudy (anonymous):

Oh wow my answer was way off! Thanks.